OpenStudy (anonymous):

find the equation of the line tangent to the curve at the point defined by the given value of t. x=sec(t),y=tan(t), at t=pi/6

hartnn (hartnn):

first find dx/dt and dy/dt can you ?

OpenStudy (anonymous):

dx/dt = sec(t)tan(t); dy/dt=sec^2(t)

hartnn (hartnn):

yes, now dy/dx is just the ratio of those, \(\huge \frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\) then just plug in t=pi/6

OpenStudy (anonymous):

dy/dx = sec^2(t)/(sec(t)tan(t)) is that right?

OpenStudy (anonymous):

dy/dx = sec(t)/tan(t)

hartnn (hartnn):

yes, thats correct. now you can either simplify that or just directly put t=pi/6

hartnn (hartnn):

Note : dy/dx will give you slope of the required line.

OpenStudy (anonymous):

ok I see! thank you!

hartnn (hartnn):

so, could you solve this Question further on your own ? if you need further help, you can just ask :) welcome ^_^

OpenStudy (anonymous):

yes, I think can solve the question now, thanks a lot!

hartnn (hartnn):

:)

hartnn (hartnn):

oh, and by the way \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)